OpenStudy (anonymous):

How many different three digit numbers greater than 240 can be formed by using three different digits from the set (1, 2, 3, 4)?

OpenStudy (anonymous):

2 in the 2 hundreds. 6 in the 3 hundreds. 6 in the 4 hundreds. So, 14 I think.

OpenStudy (anonymous):

241 243 312 314 321 324 341 342 412 413 421 423 431 432

OpenStudy (anonymous):

Assuming no repetitions allowed.

OpenStudy (ash2326):

First digit can either be 2 , 3or 4 When it's 2 the second place has to be 4 the third place can be occupied by 1 or 3 \[1\times 1\times 2=2\] When the first digit is eithere 3 or 4 We choose the first digit in 2 ways the other two digits from three no.s can be chosen as \[3\times2=6\] Total no.s =\[2\times 6=12\] Total=12+2=14

OpenStudy (anonymous):

I'm confused, some say 14, and you say 12 *_*

Directrix (directrix):

Maybe I was too "excited." Let me rethink. Go ahead and post your next question. I'll catch up.

OpenStudy (ash2326):

Saki it's 14 only

OpenStudy (anonymous):

Directrix missed the "two" numbers possible in the two hundreds, and never added those two to his/her 12. That is the difference. Right answer is 14 = 12 + 2